OpenStudy (anonymous):

the velocity of a car (in feet/second), t sec after starting from rest is given by the function f(t)= 2√t (0≤t≤30)

OpenStudy (anonymous):

find the car's position at any time t.

OpenStudy (anonymous):

\[\int_0^{30} 2t^{\frac{1}{2}}dt\]

OpenStudy (anonymous):

okay so i can pull the 2 outside and find the integral of t^(1/2)

OpenStudy (anonymous):

\[\left|2\frac{t^{\frac{1}{2}+1}}{\frac{1}{2}+1}\right|_0^{30}\]

OpenStudy (anonymous):

Yea, Exactly

OpenStudy (anonymous):

i got 2[(2/3)t^(3/2)] from 0 to 30

OpenStudy (anonymous):

yep, you got it right

OpenStudy (anonymous):